Combustion of an unknown compound containing only carbon and hydrogen produces 54.9 g of CO₂ and 45.1 g of H₂O. What is the empirical formula of the compound?
Answer #1

No. of moles of CO2 = (54.9 g) / (44.01 g/mol) = 1.25 mol

each of CO2 having 1 carbon atom in it.

So; 1.25 mol of CO2 having carbon = 1.25 mol of Carbon

No. of moles of H2O = (45.1 g) / (18.02 g/mol) = 2.502 mol

Each of H2O having No. of Hydrogens = 2

So; 2.502 mol of H2O having Hydrogens = 2 x 2.502 mol = 5.0 mol of Hydrogen

Simplest mol ratio

Carbon = (1.25 mol) / (1.25 mol) = 1

Hydrogen = (5.0 mol) / (1.25 mol) = 4

So;

the Empirical formula is: CH4